OpenStudy (chrisneedshelp):

If the first equation is multiplied by 3 and then the equations are added, the result is _____. 3x + y = 3 x - 3y = -2 A:10x = 11 B:10x = 7 C:8x = 7

OpenStudy (mathmale):

Chris, I need help myself, in understanding where you want your helpers to focus. You've typed in five equations. Are we to add them all together? What's your goal?

OpenStudy (anonymous):

Its the choices actually

OpenStudy (chrisneedshelp):

the last three are choices, sorry

OpenStudy (mathmale):

I should have seen that myself. Sorry here too. This goes to show how important it is to be clear when you type in problem and possible answers; perhaps you could label the answers a, b, c, d.

OpenStudy (unklerhaukus):

the first equation is 3x + y = 3 multiply the equation by three, mean multiply both sides by three, like this 3*(3x + y) = 3*(3) simplify this

OpenStudy (mathmale):

Following Unkle's suggestion, Type 9x + 3y = 9 followed by the 2nd equation: x - 3y = -2 ---------- Add these 2 equations together, column-by-column.